### v1

# FourierKAN-GCF: Fourier Kolmogorov-Arnold Network - An Effective and Efficient Feature Transformation for Graph Collaborative Filtering

Abstract.
Graph Collaborative Filtering (GCF) has achieved state-of-the-art performance for recommendation tasks. However, most GCF structures simplify the feature transformation and nonlinear operation during message passing in the graph convolution network (GCN). We revisit these two components and discover that a part of feature transformation and nonlinear operation during message passing in GCN can improve the representation of GCF, but increase the difficulty of training.
In this work, we propose a simple and effective graph-based recommendation model called FourierKAN-GCF. Specifically, it utilizes a novel Fourier Kolmogorov-Arnold Network (KAN) to replace the multilayer perceptron (MLP) as a part of the feature transformation during message passing in GCN, which improves the representation power of GCF and is easy to train. We further employ message dropout and node dropout strategies to improve the representation power and robustness of the model. Extensive experiments on two public datasets demonstrate the superiority of FourierKAN-GCF over most state-of-the-art methods. The implementation code is available at https://github.com/Jinfeng-Xu/FKAN-GCF.
Recommendation, Collaborative Filtering, Graph Neural Network, Kolmogorov-Arnold Network, Fourier Transform
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems;
1. Introduction
Recommender system has been widely used to alleviate information overload on the web (Xu et al., 2024). The aim of recommender system is to predict whether a user will interact with an unobserved item. To this end, collaborative filtering (CF) addresses it by learning user preferences from similar users. Many works (He et al., 2020; Wang et al., 2019; Jiang et al., 2023) harnessing graph-based recommendation to capture latent user preferences from user-item interactions have gained notable successes. The main reason is that the interactive data between users and items are naturally graph-structured. For example, NGCF (Wang et al., 2019) utilizes whole standard GCN in recommender systems, which retains the feature transformation and nonlinear operation. However, LightGCN (He et al., 2020) states that both feature transformation and nonlinear operation are unnecessary for feature extraction in recommendation field and further proposes a light GCN. Most of the subsequent graph-based models (Zhou et al., 2023a; Wei et al., 2019; Yu et al., 2023) are based on LightGCN for further exploration.
However, there are many unfair analyses of NGCF in LightGCN: a) There are two different feature transformations in NGCF, but LightGCN removed them without performing separate fine-grained ablation experiments. b) LightGCN also removed the interaction information representation part of NGCF without giving reasons. Therefore, we question: ‘Is feature transformation during message passing in GCN really unnecessary in recommendation?’
We provide an empirical analysis for NGCF and LightGCN in Section 2. Then, we state that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN, but the training process becomes more complex. In this work, we introduce a simple yet powerful graph-based recommendation model called FourierKAN-GCF. Particularly, FourierKAN-GCF incorporates a unique Fourier Kolmogorov-Arnold Network (KAN) in place of the traditional multilayer perceptron (MLP) within the feature transformation during message passing in GCNs. This substitution enhances the representational capabilities and training ease of the GCF. Additionally, we implement various dropout techniques to further boost the model’s representational strength and robustness. Extensive experiments on public datasets demonstrate the superiority of our FourierKAN-GCF over state-of-the-art methods.
2. Preliminary
2.1. NGCF Brief
NGCF retains feature transformation and nonlinear operation during the message passing in GCN, formally:
(1) $\mathbf{e}_{u}^{(l+1)}$ $=\sigma(\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}+\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}+\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
$\mathbf{e}_{i}^{(l+1)}$ $=\sigma(\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}+\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}+\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are two trainable weight matrices to perform feature transformation in each layer. $\sigma$ is the nonlinear activation function. It is worth noting that $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ can be regarded as the aggregated representation from neighbors, while $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ and $\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})$ can be regarded as the aggregated representation from interaction information. The final embeddings are calculated as:
(2) $\mathbf{\hat{e}}_{u}=\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)},\quad\mathbf{\hat{e}}_{i}=\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)},$
where ||||| | is the concatenation operation. NGCF only concatenates layer-$1$ to layer-$L$ and ignores ego layer-00, since ego layer-00 has already been considered in the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ of message passing and propagation.
2.2. LightGCN Brief
LightGCN provides an analysis of the feature transformation and nonlinear operation in NGCF. It mainly provides four observations:
1. Removing the entire feature transformation including both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ leads to consistent improvements over NGCF.
2. Removing only nonlinear operation $\sigma$ will lead to a small deterioration of performance.
3. Removing both entire feature transformation and nonlinear operation can improve performance significantly.
4. The deterioration of NGCF stems from the training difficulty rather than over-fitting.
To this end, LightGCN removes feature transformation and nonlinear operation during the message passing in GCN. Formally, LightGCN leverages the user-item graph to propagate embeddings as:
(3) $\mathbf{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{e}_{i}^{(l)}}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}},\quad\mathbf{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{e}_{u}^{(l)}}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}},$
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. LightGCN only retains the aggregation of neighbors’ representation but removes the aggregation of interaction information representation. The final embeddings are calculated as:
(4) $\mathbf{\hat{e}}_{u}=\sum_{l=0}^{L}\frac{\mathbf{e}_{u}^{(l)}}{L+1},\quad\mathbf{\hat{e}}_{i}=\sum_{l=0}^{L}\frac{\mathbf{e}_{i}^{(l)}}{L+1},$
where $L$ is the total layer number. LightGCN considers the ego layer-00 cause it removes the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ in NGCF.
We argue that observation 4 is the main reason why LightGCN performs better than NGCF in most cases. Observations 1-3 were not sufficient to verify that the entire feature transformation and nonlinear operation do not contribute to feature extraction. Besides, we argue that LightGCN verified the feature transformation not contributing to learning better features by comparing NGCF with removed both $\mathbf{W}_{1}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$, which is not a fair comparison. We point out that $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ naturally contain information about user preferences and item properties that can be adequately described through multiple feature dimensions. Therefore, $\mathbf{W}_{1}$ is an unnecessary feature transformation part. However, we point out that the interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ contains valuable interaction information, which can not be easily extracted by the heuristic rule. In this case, $\mathbf{W}_{2}$ is a necessary feature transformation part that contributes to feature extraction. In the next subsection, we provide empirical evidence for these arguments.
Table 1. Performance comparison of LightGCN, NGCF, and four variants of NGCF in terms of Recall@20 and NDCG@20.
Dataset Metrics LightGCN NGCF NGCF-f1 NGCF-f2 NGCF-i NGCF-n
MOOC Recall@20 0.3307 0.3361 0.3377$\uparrow$ 0.3357$\downarrow$ 0.3301$\downarrow$ 0.3343$\downarrow$
NDCG@20 0.1811 0.1894 0.1926$\uparrow$ 0.1897$\uparrow$ 0.1824$\downarrow$ 0.1878$\downarrow$
Games Recall@20 0.0447 0.0379 0.0414$\uparrow$ 0.0388$\uparrow$ 0.0362$\downarrow$ 0.0373$\downarrow$
NDCG@20 0.0227 0.0196 0.0209$\uparrow$ 0.0192$\downarrow$ 0.0181$\downarrow$ 0.0196--
2.3. Re-Analysis for NGCF
We implement four simplified variants of NGCF:
  • NGCF-f1, which removes feature transformation matrix $\mathbf{W}_{1}$.
  • NGCF-f2, which removes feature transformation matrix $\mathbf{W}_{2}$.
  • NGCF-i, which removes the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$.
  • NGCF-n, which removes the nonlinear operation $\sigma$.
We keep all hyper-parameter settings the same as the optimal settings of NGCF. As Table 1 shows, we conclude the findings that:
1. For both MOOC and Games datasets, removing the feature transformation matrix $\mathbf{W}_{1}$ will lead to observed improvements. Therefore, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary.
2. For MOOC dataset, removing the feature transformation matrix $\mathbf{W}_{2}$ leads to a slight deterioration. However, for Games dataset, it will lead to a small improvement. We own this to that user-item interaction in MOOC dataset has higher density than in Games dataset.
3. Removing the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will lead to a performance degradation on all datasets. It shows that this part can bring valuable interaction information.
4. Nonlinear operation only brings a small positive effect.
5. NGCF outperforms LightGCN on the MOOC dataset verifying that NGCF has better representation power than LightGCN, but the training is more difficult.
Then, we can draw some conclusions. First, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary for NGCF. Besides, the feature transformation matrix $\mathbf{W}_{2}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ is beneficial for feature extraction. Last, nonlinear operation occupies only a minor influence.
3. Methodology
In this section, we detail the overall architecture of our FourierKAN-GCF, NGCF, and LightGCN in Figure 1. Our Fourier Kolmogorov-Arnold Network provides a powerful interaction representation aggregation part in NGCF to achieve a significant performance improvement and make it easy to train.
Figure 1. Overview of FourierKAN-GCF, NGCF, and LightGCN
3.1. Kolmogorov-Arnold Network (KAN)
Kolmogorov-Arnold Network (Liu et al., 2024) is a promising alternative to Multi-Layer Perceptron (MLP). MLP is inspired by the universal approximation theorem (Cybenko, 1989). KAN focuses on the Kolmogorov-Arnold representation theorem (Kolmogorov, 1961). KAN can be formed as:
(5) $\operatorname{KAN}=f(\mathbf{x})=\sum_{q=1}^{2n+1}\Phi_{q}(\sum_{p=1}^{n}\phi_{q,p}(\mathbf{x}_{p})),$
where $\phi_{q,p}$ are univariate functions that map each input variable $\mathbf{x}_{p}$ such $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. $\phi_{q,p}(\mathbf{x}_{p})$ is trainable activation function. In the KolmogovArnold theorem, the inner functions form a KAN layer with $n_{in}$ = $n$ and $n_{out}$ = $12n+$ 1, and the outer functions form a KAN layer with $n_{in}$ = $12n+$ 1 and $n_{out}$ = $n$. So the Kolmogorov-Arnold representations in this formula are simply compositions of two KAN layers. A useful trick is that it includes a basis function $b(\mathbf{x})$ such that the activation function $\phi(\mathbf{x})$ is the sum of the basis function $b(\mathbf{x})$ and the spline function:
(6) $\phi(\mathbf{x})=\mathbf{w}(b(\mathbf{x})+\operatorname{spline}(\mathbf{x})).$
$\operatorname{spline}(\mathbf{x})$ is parameterized as a linear combination of B-splines:
(7) $b(x)=\operatorname{silu}(\mathbf{x})=\frac{x}{1+e^{-\mathbf{x}}},\quad\operatorname{spline}(\mathbf{x})=\sum_{i}c_{i}B_{i}(\mathbf{x}),$
where $c_{i}$ is trainable.
3.2. Fourier KAN
However, KAN is more difficult to train than MLP due to the spline function, which does not satisfy our motivation to provide a more effective and efficient method to replace $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ part. The essence of KAN is to form an arbitrary function by superposition of multiple nonlinear functions. Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions. Naturally, the Fourier transform (Nussbaumer and Nussbaumer, 1982) is an interesting choice. Therefore, we propose the following equation:
(8) $\phi_{F}(\mathbf{x})=\sum_{i=1}^{d}\sum_{k=1}^{g}\left(\cos\left(k\mathbf{x}_{i}\right)\cdot a_{ik}+\sin\left(k\mathbf{x}_{i}\right)\cdot b_{ik}\right),$
where $d$ is the dimension number of features. Fourier coefficients $a_{ik}$ and $b_{ik}$ are trainable. Hyper-parameter $g$ is the gridsize, which plays a critical role in controlling the number of terms (frequencies) used in the Fourier series expansion. Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier transform corresponding to each input dimension. The Fourier transform has a significant advantage in computational efficiency and solves the training difficulty caused by the spline function.
3.3. FourierKAN-GCF
The message passing in our FourierKAN-GCF is defined as:
(9) $\mathbf{e}_{u}^{(l+1)}$ $=\sigma(\mathbf{e}_{u}^{(l)}+\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{e}_{i}^{(l)}+\phi_{F}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
$\mathbf{e}_{i}^{(l+1)}$ $=\sigma(\mathbf{e}_{i}^{(l)}+\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{e}_{u}^{(l)}+\phi_{F}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
where $\phi_{F}(\cdot)$ is simplified single layer Fourier KAN function. We remove the unnecessary transform matrix $\mathbf{W}_{1}$ in NGCF and utilize our Fourier KAN function to replace the transform matrix $\mathbf{W}_{2}$. The final user embedding and item embedding are calculated as:
(10) $\mathbf{\hat{e}}_{u}=\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)},\quad\mathbf{\hat{e}}_{i}=\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{i}^{(L)},$
where ||||| | is the concatenation operation.
3.4. Dropout Strategies
To further prevent the over-fitting problem for FourierKAN-GCF, we adopt message dropout and node dropout strategies. For the message dropout, we randomly change some message passing being propagated in Eq 9 by zero, with a probability $1-p_{m}$, where $p_{m}$ is the message dropout ratio. For the node dropout, we randomly drop $1-p_{n}$ nodes of the matrix, where $p_{n}$ is the node dropout ratio.
Table 2. Performance comparison of baselines and FourierKAN-GCF in terms of R@K and N@K. The superscript indicates the improvement is statistically significant where the p-value is less than 0.05.
Datasets 1MOOC 2Games
Metrics R@10 R@20 R@50 N@10 N@20 N@50 R@10 R@20 R@50 N@10 N@20 N@50
BPR-MF 0.2401 0.3353 0.4813 0.1571 0.1898 0.2261 0.0210 0.0369 0.0699 0.0135 0.0183 0.0265
BUIR 0.2354 0.3196 0.4801 0.1589 0.1835 0.2224 0.0227 0.0384 0.0749 0.0143 0.0192 0.0282
NGCF 0.2453 0.3361 0.4799 0.1722 0.1894 0.2349 0.0231 0.0379 0.0782 0.0142 0.0196 0.0274
LR-GCCF 0.2466 0.3336 0.4809 0.1678 0.1938 0.2294 0.0253 0.0440 0.0815 0.0171 0.0224 0.0317
LightGCN 0.2402 0.3307 0.4773 0.1581 0.1811 0.2217 0.0271 0.0447 0.0844 0.0174 0.0227 0.0326
UltraGCN 0.2372 0.3194 0.4701 0.1737 0.1962 0.2307 0.0277 0.0459 0.0844 0.0173 0.0230 0.0331
IMP-GCN 0.2170 0.2788 0.4183 0.1531 0.1717 0.2057 0.0273 0.0461 0.0839 0.0171 0.0232 0.0323
w/o MD 0.2518 0.3523 0.4912 0.1814 0.2116 0.2449 0.0267 0.0452 0.0825 0.0165 0.0221 0.0314
w/o ND 0.2603 0.3527 0.4839 0.1800 0.2071 0.2439 0.0257 0.0452 0.0809 0.0161 0.0219 0.0309
KAN-GCF 0.2536 0.3417 0.4984 0.1758 0.2024 0.2396 0.0272 0.0451 0.0837 0.0172 0.0229 0.0325
FourierKAN-GCF 0.2595 0.3564 0.5065 0.1866 0.2147 0.2462 0.0287 0.0473 0.0856 0.0176 0.0252 0.0342
1Dataset can be accessed at http://moocdata.cn. 2Dataset can be accessed at http://jmcauley.ucsd.edu/data/amazon/links.html.
3.5. Model Training
To learn model parameters, we adopt the Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) loss function as our optimization criterion. The core objective of BPR is to enhance the divergence in the predictive preference between positive and negative items within each user-item triplet $(u,i_{p},i_{n})\in\mathcal{D}$, where $\mathcal{D}$ signifies the collection of training data, the term positive item $p$ pertains to an item with which the user $u$ has interacted, whereas the negative item $n$ is selected randomly from the pool of items without interaction with user $u$.
(11) $\mathcal{L}=\sum_{(u,i_{p},i_{n})\in\mathcal{D}}-\ln\sigma(\mathbf{\hat{e}}_{u}^{T}\mathbf{\hat{e}}_{i_{p}}-\mathbf{\hat{e}}_{u}^{T}\mathbf{\hat{e}}_{i_{n}})+\lambda\|\mathbf{\Theta}\|^{2},$
where $\lambda$ controls the $L_{2}$ regularization strength, $\sigma$ is the Sigmoid function, and $\mathbf{\Theta}$ denotes model parameters.
Table 3. Statistics of experimental datasets.
Dataset # Users # Items # Interaction Sparsity
MOOC 82535 1302 458453 99.57%
Games 50677 16897 454529 99.95%
4. Experiment
4.1. Datasets and Evaluation Metrics
We conduct our experiments on two real-world datasets: MOOC and Anazon Video Games (Games for short). Details can be found in Table 3. For a fair comparison, we follow previous work (Zhou et al., 2023b) to sort all the observed user-item interactions of each dataset in chronological order based on the interaction timestamps. Then, we split each dataset with a ratio of 7:1:2 for training, validation, and testing. Regarding evaluation metrics, we adopt two well-established metrics: Recall@K (R@K) and NDCG@K (N@K). We report the average metrics of all users in the test dataset among $K$ = 10, $K$ = 20, and $K$ = 50.
4.2. Baselines and Experimental Settings
To verify the effectiveness of FourierKAN-GCF, we select two traditional matrix factorization models (BPR-MF (Rendle et al., 2012) and BUIR (Lee et al., 2021)) and five state-of-the-art GCN-based models (NGCF (Wang et al., 2019), LR-GCCF (Chen et al., 2020), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), and IMP-GCN (Liu et al., 2021)) for comparison. For a fair comparison, we implement all baselines under a unified framework (Zhou et al., 2023b), fix the embedding size of both users and items to 64 for all models, initialize the embedding parameters with the Xavier initialization (Glorot and Bengio, 2010), and use Adam (Kingma and Ba, 2014) as the optimizer. Besides, we carefully tune the hyper-parameters of each baseline following their published papers to find optimal settings. For our FourierKAN-GCF, we tune the $\lambda$ of $L_{2}$ regularization term in {0, 1e-2, 1e-1, 1, 10}, and tune the layer number $L$ from 1 to 4. The gridsize $g$ of Fourier KAN is searched from {1, 2, 4, 8}. We further grid search the message dropout ratio $p_{m}$ and node dropout ratio $p_{n}$ in {0.0, 0.1, 0.2, 0.3}.
4.3. Performance Comparison
The results of our experiments are listed in Table 2. We can identify that the majority of GCN-based methods outperform traditional matrix factorization methods, which is a common trend. Our FourierKAN-GCF outperforms all baselines in two datasets across various metrics. We owe our superiority to FourierKAN, which is easier to train and has greater representation power than MLP. It is worth noting that FourierKAN-GCF can adjust the training difficulty by adjusting the grid size $g$.
4.4. Ablation Study
Table 2 also demonstrates the significance of dropout strategies. We use w/o MD and w/o ND to denote without message dropout and without node dropout, respectively. This ablation study shows that both message dropout and node dropout play a distinct role in improving model representation power and model robustness. We use KAN-GCF to denote a variant that utilizes our standard KAN function to replace transform matrix $\mathbf{W}_{2}$ in NGCF. It shows that standard KAN is better than MLP but worse than our Fourier KAN.
4.5. Sensitivity Analysis
To analyze the hyper-parameter sensitivity of FourierKAN-GCF, we test the performance of FourierKAN-GCF on two datasets with different hyper-parameters. The optimal layer number $L$ is 3 on all datasets, shown in Figure LABEL:fig:l. Besides, Figure LABEL:fig:g demonstrates that for MOOC and Games datasets, optimal ($p_{m}$, $p_{n}$) pairs are (0.2, 0.2) and (0.1, 0.1), respectively. As Figure LABEL:fig:dropout illustrated, for the relatively dense dataset MOOC, $g$ = 4 is the best grid size. For the relatively sparse dataset Games, $g$ = 2 is the best grid size. This further demonstrates that FourierKAN can be easily adapted to various datasets.
5. Conclusion
In this paper, we revisit the processes of feature transformation and nonlinear operations during message passing in the NGCF. We assert that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN but add complexity in training. To reduce the training difficulty while benefiting from these representational enhancements, we introduce a novel feature transformation method called FourierKAN. Inspired by KAN, FourierKAN employs a Fourier transform instead of the Spline function in standard KAN, which further reduces the difficulty of training while increasing the representation power. Furthermore, based on FourierKAN, we introduce a simple and effective graph-based recommendation model (FourierKAN-GCF) with extra dropout strategies. Extensive experiments on public datasets demonstrate the superiority of our model over the state-of-the-art methods.

### v2

@@
 The implementation code is available at https://github.com/Jinfeng-Xu/FKAN-GCF.
-Recommendation, Collaborative Filtering, Graph Neural Network, Kolmogorov-Arnold Network, Fourier Transform
+Recommendation, Collaborative Filtering, Graph Neural Network, Kolmogorov-Arnold Network, Fourier Coefficients
 copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems;
@@
 Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions.
-Naturally, the Fourier transform (Nussbaumer and Nussbaumer, 1982) is an interesting choice.
+Naturally, the Fourier Coefficients (Nussbaumer and Nussbaumer, 1982) is an interesting choice.
 Therefore, we propose the following equation:
@@
 Hyper-parameter $g$ is the gridsize, which plays a critical role in controlling the number of terms (frequencies) used in the Fourier series expansion.
-Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier transform corresponding to each input dimension.
-The Fourier transform has a significant advantage in computational efficiency and solves the training difficulty caused by the spline function.
+Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier Coefficients corresponding to each input dimension.
+The Fourier Coefficients has a significant advantage in computational efficiency and solves the training difficulty caused by the spline function.
 3.3. FourierKAN-GCF
@@
 To reduce the training difficulty while benefiting from these representational enhancements, we introduce a novel feature transformation method called FourierKAN.
-Inspired by KAN, FourierKAN employs a Fourier transform instead of the Spline function in standard KAN, which further reduces the difficulty of training while increasing the representation power.
+Inspired by KAN, FourierKAN employs a Fourier Coefficients instead of the Spline function in standard KAN, which further reduces the difficulty of training while increasing the representation power.
 Furthermore, based on FourierKAN, we introduce a simple and effective graph-based recommendation model (FourierKAN-GCF) with extra dropout strategies.

### v3

@@
-FourierKAN-GCF: Fourier Kolmogorov-Arnold Network - An Effective and Efficient Feature Transformation for Graph Collaborative Filtering
+Enhancing Graph Collaborative Filtering with FourierKAN Feature Transformation
 Abstract.
-Graph Collaborative Filtering (GCF) has achieved state-of-the-art performance for recommendation tasks.
-However, most GCF structures simplify the feature transformation and nonlinear operation during message passing in the graph convolution network (GCN).
-We revisit these two components and discover that a part of feature transformation and nonlinear operation during message passing in GCN can improve the representation of GCF, but increase the difficulty of training.
-In this work, we propose a simple and effective graph-based recommendation model called FourierKAN-GCF.
-Specifically, it utilizes a novel Fourier Kolmogorov-Arnold Network (KAN) to replace the multilayer perceptron (MLP) as a part of the feature transformation during message passing in GCN, which improves the representation power of GCF and is easy to train.
-We further employ message dropout and node dropout strategies to improve the representation power and robustness of the model.
-Extensive experiments on two public datasets demonstrate the superiority of FourierKAN-GCF over most state-of-the-art methods.
-The implementation code is available at https://github.com/Jinfeng-Xu/FKAN-GCF.
+Graph Collaborative Filtering (GCF) has emerged as a dominant paradigm in modern recommendation systems, excelling at modeling complex user-item interactions and capturing high-order collaborative signals through graph-structured learning.
+Most existing GCF models predominantly rely on simplified graph architectures like LightGCN, which strategically remove feature transformation and activation functions from vanilla graph convolution networks.
+Through systematic analysis, we reveal that feature transformation in message propagation can enhance model representation, though at the cost of increased training difficulty.
+To this end, we propose FourierKAN-GCF, a novel GCN framework that adopts Fourier Kolmogorov-Arnold Networks as efficient transformation modules within graph propagation layers.
+This design enhances model representation while decreasing training difficulty.
+Our FourierKAN-GCF can achieve higher recommendation performance than most widely used GCF backbone models.
+In addition, it can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance.
+Extensive experiments on three public datasets demonstrate the superiority of FourierKAN-GCF.
 Recommendation, Collaborative Filtering, Graph Neural Network, Kolmogorov-Arnold Network, Fourier Coefficients
-copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems;
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3760909isbn: 979-8-4007-2040-6/2025/11
 1. Introduction
-Recommender system has been widely used to alleviate information overload on the web (Xu et al., 2024).
-The aim of recommender system is to predict whether a user will interact with an unobserved item.
-To this end, collaborative filtering (CF) addresses it by learning user preferences from similar users.
-Many works (He et al., 2020; Wang et al., 2019; Jiang et al., 2023) harnessing graph-based recommendation to capture latent user preferences from user-item interactions have gained notable successes.
-The main reason is that the interactive data between users and items are naturally graph-structured.
-For example, NGCF (Wang et al., 2019) utilizes whole standard GCN in recommender systems, which retains the feature transformation and nonlinear operation.
-However, LightGCN (He et al., 2020) states that both feature transformation and nonlinear operation are unnecessary for feature extraction in recommendation field and further proposes a light GCN.
-Most of the subsequent graph-based models (Zhou et al., 2023a; Wei et al., 2019; Yu et al., 2023) are based on LightGCN for further exploration.
+Recommender systems are widely used to alleviate information overload on the Web (Chen et al., 2020; He et al., 2017; Chen et al., 2025a, b; Xu et al., 2025e, a), aiming to recommend suitable items for users based on their historical behavior.
+Collaborative filtering (CF) addresses this by learning user preferences from similar users.
+Recently, graph-based CF (GCF) models (He et al., 2020; Wang et al., 2019; Jiang et al., 2023; Xu et al., 2025d, c, b, 2024b) have achieved notable success, as user-item interactions are naturally graph-structured.
+For example, NGCF (Wang et al., 2019) adopts whole standard GCN in recommender systems, which retains the feature transformation and nonlinear operation.
+However, LightGCN (He et al., 2020) states that both feature transformation and nonlinear operation are unnecessary in the recommendation field and further proposes a lightweight GCN.
+Consequently, most subsequent GCF models (Chen et al., 2020; Cai et al., 2023; Xu et al., 2024a) adopt LightGCN as the backbone for further exploration.
 However, there are many unfair analyses of NGCF in LightGCN: a) There are two different feature transformations in NGCF, but LightGCN removed them without performing separate fine-grained ablation experiments.
@@
 We provide an empirical analysis for NGCF and LightGCN in Section 2.
-Then, we state that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN, but the training process becomes more complex.
+Then, we state that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN, but increases the training difficulty.
 In this work, we introduce a simple yet powerful graph-based recommendation model called FourierKAN-GCF.
-Particularly, FourierKAN-GCF incorporates a unique Fourier Kolmogorov-Arnold Network (KAN) in place of the traditional multilayer perceptron (MLP) within the feature transformation during message passing in GCNs.
-This substitution enhances the representational capabilities and training ease of the GCF.
-Additionally, we implement various dropout techniques to further boost the model’s representational strength and robustness.
-Extensive experiments on public datasets demonstrate the superiority of our FourierKAN-GCF over state-of-the-art methods.
+Specifically, FourierKAN-GCF incorporates a unique Fourier Kolmogorov-Arnold Network (KAN) in place of the traditional multilayer perceptron (MLP) within the feature transformation during message passing in GCNs.
+This substitution enhances the representational capabilities and reduces the difficulty of training for GCFs.
+FourierKAN-GCF can achieve higher recommendation performance than most widely-used GCF backbone models.
+In addition, FourierKAN-GCF can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance.
+Extensive experiments on public datasets demonstrate the superiority of FourierKAN-GCF over state-of-the-art methods.
+Our work is intended to reawaken researchers’ thinking about feature transformation in GCF, rather than arbitrarily removing it.
 2. Preliminary
 2.1. NGCF Brief
-NGCF retains feature transformation and nonlinear operation during the message passing in GCN, formally:
+NGCF (Wang et al., 2019) retains feature transformation and nonlinear operation during the message passing in GCN, formally:
 (1) $\mathbf{e}_{u}^{(l+1)}$ $=\sigma(\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}+\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}+\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
@@
 It is worth noting that $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ can be regarded as the aggregated representation from neighbors, while $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ and $\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})$ can be regarded as the aggregated representation from interaction information.
-The final embeddings are calculated as:
-(2) $\mathbf{\hat{e}}_{u}=\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)},\quad\mathbf{\hat{e}}_{i}=\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)},$
-where ||||| | is the concatenation operation.
-NGCF only concatenates layer-$1$ to layer-$L$ and ignores ego layer-00, since ego layer-00 has already been considered in the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ of message passing and propagation.
+The final embeddings are calculated by $\mathbf{\hat{e}}_{u}=\{\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)}\}$ and $\mathbf{\hat{e}}_{i}=\{\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)}\}$, where $||$ is the concatenation operation.
+NGCF only concatenates layer-$1$ to layer-$L$ and ignores ego layer-0, since ego layer-0 has already been considered in the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ of message passing and propagation.
 2.2. LightGCN Brief
-LightGCN provides an analysis of the feature transformation and nonlinear operation in NGCF. It mainly provides four observations:
-1. Removing the entire feature transformation including both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ leads to consistent improvements over NGCF.
-2. Removing only nonlinear operation $\sigma$ will lead to a small deterioration of performance.
-3. Removing both entire feature transformation and nonlinear operation can improve performance significantly.
-4. The deterioration of NGCF stems from the training difficulty rather than over-fitting.
-To this end, LightGCN removes feature transformation and nonlinear operation during the message passing in GCN. Formally, LightGCN leverages the user-item graph to propagate embeddings as:
+LightGCN (He et al., 2020) analyzes the feature transformation and nonlinear operation in NGCF.
+It offers four observations: a) Removing the entire feature transformation, including $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, leads to consistent improvements over NGCF.
+b) Removing only nonlinear operation $\sigma$ will lead to a small deterioration of performance.
+c) Removing both entire feature transformation and nonlinear operation can improve performance significantly.
+d) The deterioration of NGCF stems from the training difficulty rather than over-fitting.
+To this end, LightGCN removes feature transformation and nonlinear operation during the message passing in GCN. Formally, the user-item graph to propagate embeddings as:
 (2) $\mathbf{e}_{u}^{(l+1)}=\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{e}_{i}^{(l)}}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}},\quad\mathbf{e}_{i}^{(l+1)}=\sum_{u\in\mathcal{N}_{i}}\frac{\mathbf{e}_{u}^{(l)}}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}},$
@@
 LightGCN only retains the aggregation of neighbors’ representation but removes the aggregation of interaction information representation.
-The final embeddings are calculated as:
-(4) $\mathbf{\hat{e}}_{u}=\sum_{l=0}^{L}\frac{\mathbf{e}_{u}^{(l)}}{L+1},\quad\mathbf{\hat{e}}_{i}=\sum_{l=0}^{L}\frac{\mathbf{e}_{i}^{(l)}}{L+1},$
-where $L$ is the total layer number. LightGCN considers the ego layer-00 cause it removes the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ in NGCF.
-We argue that observation 4 is the main reason why LightGCN performs better than NGCF in most cases.
-Observations 1-3 were not sufficient to verify that the entire feature transformation and nonlinear operation do not contribute to feature extraction.
+The final embeddings are calculated as $\mathbf{\hat{e}}_{u}=\sum_{l=0}^{L}\frac{1}{L+1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{\hat{e}}_{i}=\sum_{l=0}^{L}\frac{1}{L+1}\mathbf{e}_{i}^{(l)}$, where $L$ is the total layer number.
+LightGCN considers the ego layer-0 cause it removes the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$.
+We argue that observation (d) is the main reason why LightGCN performs better than NGCF in most cases.
+Observations (a-c) were not sufficient to verify that the entire feature transformation and nonlinear operation do not contribute to feature extraction.
 Besides, we argue that LightGCN verified the feature transformation not contributing to learning better features by comparing NGCF with removed both $\mathbf{W}_{1}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$, which is not a fair comparison.
@@
 In this case, $\mathbf{W}_{2}$ is a necessary feature transformation part that contributes to feature extraction.
-In the next subsection, we provide empirical evidence for these arguments.
-Table 1. Performance comparison of LightGCN, NGCF, and four variants of NGCF in terms of Recall@20 and NDCG@20.
-Dataset Metrics LightGCN NGCF NGCF-f1 NGCF-f2 NGCF-i NGCF-n
-MOOC Recall@20 0.3307 0.3361 0.3377$\uparrow$ 0.3357$\downarrow$ 0.3301$\downarrow$ 0.3343$\downarrow$
-NDCG@20 0.1811 0.1894 0.1926$\uparrow$ 0.1897$\uparrow$ 0.1824$\downarrow$ 0.1878$\downarrow$
-Games Recall@20 0.0447 0.0379 0.0414$\uparrow$ 0.0388$\uparrow$ 0.0362$\downarrow$ 0.0373$\downarrow$
-NDCG@20 0.0227 0.0196 0.0209$\uparrow$ 0.0192$\downarrow$ 0.0181$\downarrow$ 0.0196--
+In the next subsection, we provide an empirical analysis.
+Table 1. Performance comparison of LightGCN, NGCF, and six variants of NGCF in terms of Recall@20 (R@20) and NDCG@20 (N@20).
+Datasets MOOC Amazon Gowalla
+Metrics R@20 N@20 R@20 N@20 R@20 N@20
+LightGCN 0.3307 0.1811 0.0447 0.0227 0.1830 0.1152
+NGCF 0.3361 0.1894 0.0379 0.0196 0.1755 0.1013
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
 2.3. Re-Analysis for NGCF
-We implement four simplified variants of NGCF:
-  • NGCF-f1, which removes feature transformation matrix $\mathbf{W}_{1}$.
-  • NGCF-f2, which removes feature transformation matrix $\mathbf{W}_{2}$.
-  • NGCF-i, which removes the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$.
-  • NGCF-n, which removes the nonlinear operation $\sigma$.
-We keep all hyper-parameter settings the same as the optimal settings of NGCF. As Table 1 shows, we conclude the findings that:
-1.
-For both MOOC and Games datasets, removing the feature transformation matrix $\mathbf{W}_{1}$ will lead to observed improvements.
+We implement six variants of NGCF: 1) NGCF-f1 removes feature transformation matrix $\mathbf{W}_{1}$.
+2) NGCF-f2 removes feature transformation matrix $\mathbf{W}_{2}$.
+3) NGCF-i removes the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$.
+4) NGCF-f1-f2 removes both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$.
+5) NGCF-f1-i removes both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$.
+6) NGCF-n removes the nonlinear operation $\sigma$.
+We keep all optimal hyper-parameter settings as the NGCF reported.
+As Table 1 shows, we conclude the findings: a) For all three datasets, removing the feature transformation matrix $\mathbf{W}_{1}$ will lead to observed improvements.
 Therefore, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary.
-2.
-For MOOC dataset, removing the feature transformation matrix $\mathbf{W}_{2}$ leads to a slight deterioration.
-However, for Games dataset, it will lead to a small improvement.
-We own this to that user-item interaction in MOOC dataset has higher density than in Games dataset.
-3.
-Removing the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will lead to a performance degradation on all datasets.
-It shows that this part can bring valuable interaction information.
-4. Nonlinear operation only brings a small positive effect.
-5. NGCF outperforms LightGCN on the MOOC dataset verifying that NGCF has better representation power than LightGCN, but the training is more difficult.
+b) For the MOOC dataset, removing the feature transformation matrix $\mathbf{W}_{2}$ leads to a slight deterioration.
+However, for the other two datasets, it will lead to a small improvement.
+Moreover, removing both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ will lead to a small deterioration than only removing $\mathbf{W}_{1}$ for all datasets.
+We own this phenomenon to that removing only $\mathbf{W}_{1}$ can make the model optimization focus on training $\mathbf{W}_{2}$, and remove both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ will lose the representation power from feature transformation.
+c) Removing the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will lead to a performance degradation on all datasets.
+Additionally, removing both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will make a obviously degration than only remove $\mathbf{W}_{1}$ and even NGCF.
+This shows that this part can bring valuable information about interaction.
+d) Nonlinear operation only brings a small positive effect.
+e) NGCF outperforms LightGCN on the MOOC dataset verifying that NGCF has better representation power than LightGCN, but the training is more difficult (He et al., 2020).
 Then, we can draw some conclusions.
@@
 Besides, the feature transformation matrix $\mathbf{W}_{2}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ is beneficial for feature extraction.
-Last, nonlinear operation occupies only a minor influence.
+Last, nonlinear operation only has a minor influence.
+To this end, a feature transformation approach, simpler to train than MLPs but with strong representational capabilities, could be effective in graph learning for recommendations.
+KAN (Liu et al., 2024) is regarded as a promising alternative to MLP.
+While KAN shares the same theoretically unlimited representational capacity as MLP (Wang et al., 2024), the practical representational capacity of MLP is constrained by the hidden dimensionality.
+In contrast, the practical representational capacity of KAN depends on its ability to fit trainable activation functions.
+Our work can be seen as exploring adopting KAN as a feature transformation approach within GCF.
 3. Methodology
-In this section, we detail the overall architecture of our FourierKAN-GCF, NGCF, and LightGCN in Figure 1.
-Our Fourier Kolmogorov-Arnold Network provides a powerful interaction representation aggregation part in NGCF to achieve a significant performance improvement and make it easy to train.
-Figure 1. Overview of FourierKAN-GCF, NGCF, and LightGCN
+We detail the overview of FourierKAN-GCF [^1] Code is available at: https://anonymous.4open.science/r/FourierKAN-GCF-r.
+in Figure 1.
+FourierKAN-GCF introduces a promising feature transformation for GCF, which significantly boosts performance and simplifies training.
+Note that FourierKAN-GCF can be adopted as the backbone in existing advanced self-supervised models.
+Figure 1. Overview of FourierKAN-GCF.
 3.1. Kolmogorov-Arnold Network (KAN)
@@
 KAN focuses on the Kolmogorov-Arnold representation theorem (Kolmogorov, 1961).
-KAN can be formed as:
+Specifically, unlike MLPs, which have fixed activation functions on nodes, KANs contain learnable activation functions on edges (weights).
+This unique architecture enables KANs to learn nonlinear functions more effectively.
+Formally:
 (3) $\operatorname{KAN}=f(\mathbf{x})=\sum_{q=1}^{2n+1}\Phi_{q}(\sum_{p=1}^{n}\phi_{q,p}(\mathbf{x}_{p})),$
@@
 $\phi_{q,p}(\mathbf{x}_{p})$ is trainable activation function.
-In the KolmogovArnold theorem, the inner functions form a KAN layer with $n_{in}$ = $n$ and $n_{out}$ = $12n+$ 1, and the outer functions form a KAN layer with $n_{in}$ = $12n+$ 1 and $n_{out}$ = $n$.
+In the KolmogovArnold theorem, the inner functions form a KAN layer with $_{in} = nn$ and $n_{out}$ = $2n+1$, and the outer functions form a KAN layer with $n_{in}$ = $2n+1$ and $_{out} = nn$.
 So the Kolmogorov-Arnold representations in this formula are simply compositions of two KAN layers.
-A useful trick is that it includes a basis function $b(\mathbf{x})$ such that the activation function $\phi(\mathbf{x})$ is the sum of the basis function $b(\mathbf{x})$ and the spline function:
-(6) $\phi(\mathbf{x})=\mathbf{w}(b(\mathbf{x})+\operatorname{spline}(\mathbf{x})).$
-$\operatorname{spline}(\mathbf{x})$ is parameterized as a linear combination of B-splines:
-(7) $b(x)=\operatorname{silu}(\mathbf{x})=\frac{x}{1+e^{-\mathbf{x}}},\quad\operatorname{spline}(\mathbf{x})=\sum_{i}c_{i}B_{i}(\mathbf{x}),$
-where $c_{i}$ is trainable.
+A useful trick is that it includes a basis function $b(\mathbf{x})$ such that the activation function $\operatorname{silu}(\mathbf{x})=\frac{x}{1+e^{-\mathbf{x}}}$.
+$\phi(\mathbf{x})=\mathbf{w}(b(\mathbf{x})+\operatorname{spline}(\mathbf{x}))$ is the sum of the basis function $b(\mathbf{x})$ and function $\operatorname{spline}(\mathbf{x})=\sum_{i}c_{i}B_{i}(\mathbf{x})$ is a linear combination of B-splines, where $c_{i}$ is a trainable parameter.
 3.2. Fourier KAN
-However, KAN is more difficult to train than MLP due to the spline function, which does not satisfy our motivation to provide a more effective and efficient method to replace $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ part.
-The essence of KAN is to form an arbitrary function by superposition of multiple nonlinear functions.
+To further reduce the training difficulty and be able to adopt to different scenarios.
 Our goal can be converted into finding the split from a complex function into multiple relatively simple nonlinear functions.
-Naturally, the Fourier Coefficients (Nussbaumer and Nussbaumer, 1982) is an interesting choice.
+Naturally, the Fourier Coefficients (Nussbaumer and Nussbaumer, 1982) is a potential choice.
 Therefore, we propose the following equation:
@@
 Specifically, $g$ determines how many different sine and cosine terms are included in the Fourier Coefficients corresponding to each input dimension.
-The Fourier Coefficients has a significant advantage in computational efficiency and solves the training difficulty caused by the spline function.
+The Fourier Coefficients has a significant advantage in computational efficiency and reduces the training difficulty caused by the spline function.
 3.3. FourierKAN-GCF
-The message passing in our FourierKAN-GCF is defined as:
+The message passing in FourierKAN-GCF is defined as:
 (5) $\mathbf{e}_{u}^{(l+1)}$ $=\sigma(\mathbf{e}_{u}^{(l)}+\sum_{i\in\mathcal{N}_{u}}\frac{\mathbf{e}_{i}^{(l)}+\phi_{F}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})}{\sqrt{|\mathcal{N}_{u}||\mathcal{N}_{i}|}}),$
@@
 We remove the unnecessary transform matrix $\mathbf{W}_{1}$ in NGCF and utilize our Fourier KAN function to replace the transform matrix $\mathbf{W}_{2}$.
-The final user embedding and item embedding are calculated as:
-(10) $\mathbf{\hat{e}}_{u}=\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)},\quad\mathbf{\hat{e}}_{i}=\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{i}^{(L)},$
-where ||||| | is the concatenation operation.
+The final user embedding and item embedding are calculated by $\mathbf{\hat{e}}_{u}=\{\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)}\}$ and $\mathbf{\hat{e}}_{i}=\{\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)}\}$, where $||$ is the concatenation operation.
 3.4. Dropout Strategies
-To further prevent the over-fitting problem for FourierKAN-GCF, we adopt message dropout and node dropout strategies.
-For the message dropout, we randomly change some message passing being propagated in Eq 9 by zero, with a probability $1-p_{m}$, where $p_{m}$ is the message dropout ratio.
-For the node dropout, we randomly drop $1-p_{n}$ nodes of the matrix, where $p_{n}$ is the node dropout ratio.
-Table 2. Performance comparison of baselines and FourierKAN-GCF in terms of R@K and N@K. The superscript indicates the improvement is statistically significant where the p-value is less than 0.05.
-Datasets 1MOOC 2Games
-Metrics R@10 R@20 R@50 N@10 N@20 N@50 R@10 R@20 R@50 N@10 N@20 N@50
-BPR-MF 0.2401 0.3353 0.4813 0.1571 0.1898 0.2261 0.0210 0.0369 0.0699 0.0135 0.0183 0.0265
-BUIR 0.2354 0.3196 0.4801 0.1589 0.1835 0.2224 0.0227 0.0384 0.0749 0.0143 0.0192 0.0282
-NGCF 0.2453 0.3361 0.4799 0.1722 0.1894 0.2349 0.0231 0.0379 0.0782 0.0142 0.0196 0.0274
-LR-GCCF 0.2466 0.3336 0.4809 0.1678 0.1938 0.2294 0.0253 0.0440 0.0815 0.0171 0.0224 0.0317
-LightGCN 0.2402 0.3307 0.4773 0.1581 0.1811 0.2217 0.0271 0.0447 0.0844 0.0174 0.0227 0.0326
-UltraGCN 0.2372 0.3194 0.4701 0.1737 0.1962 0.2307 0.0277 0.0459 0.0844 0.0173 0.0230 0.0331
-IMP-GCN 0.2170 0.2788 0.4183 0.1531 0.1717 0.2057 0.0273 0.0461 0.0839 0.0171 0.0232 0.0323
-w/o MD 0.2518 0.3523 0.4912 0.1814 0.2116 0.2449 0.0267 0.0452 0.0825 0.0165 0.0221 0.0314
-w/o ND 0.2603 0.3527 0.4839 0.1800 0.2071 0.2439 0.0257 0.0452 0.0809 0.0161 0.0219 0.0309
-KAN-GCF 0.2536 0.3417 0.4984 0.1758 0.2024 0.2396 0.0272 0.0451 0.0837 0.0172 0.0229 0.0325
-FourierKAN-GCF 0.2595 0.3564 0.5065 0.1866 0.2147 0.2462 0.0287 0.0473 0.0856 0.0176 0.0252 0.0342
-1Dataset can be accessed at http://moocdata.cn. 2Dataset can be accessed at http://jmcauley.ucsd.edu/data/amazon/links.html.
+To mitigate overfitting in FourierKAN-GCF, we employ message dropout and node dropout strategies, similar to NGCF (Wang et al., 2019).
+In message dropout, a fraction $1-p_{m}$ of message passing in Eq.
+5 is randomly set to zero, where $p_{m}$ is the dropout ratio.
+In node dropout, $1-p_{n}$ of the nodes in the matrix are randomly dropped, with $p_{n}$ as the dropout ratio.
 3.5. Model Training
-To learn model parameters, we adopt the Bayesian Personalized Ranking (BPR) (Rendle et al., 2012) loss function as our optimization criterion.
+For model optimization, we adopt the Bayesian Personalized Ranking (BPR) (Rendle et al., 2009) loss function as our optimization criterion.
 The core objective of BPR is to enhance the divergence in the predictive preference between positive and negative items within each user-item triplet $(u,i_{p},i_{n})\in\mathcal{D}$, where $\mathcal{D}$ signifies the collection of training data, the term positive item $p$ pertains to an item with which the user $u$ has interacted, whereas the negative item $n$ is selected randomly from the pool of items without interaction with user $u$.
@@
 where $\lambda$ controls the $L_{2}$ regularization strength, $\sigma$ is the Sigmoid function, and $\mathbf{\Theta}$ denotes model parameters.
+4. Experiment
+In this section, we compare the performance of FourierKAN-GCF with popular graph-based backbone models and demonstrate its compatibility with advanced self-supervised graph-based models.
 Table 2. Statistics of experimental datasets.
 Dataset # Users # Items # Interaction Sparsity
-MOOC 82535 1302 458453 99.57%
-Games 50677 16897 454529 99.95%
-4. Experiment
+MOOC 82,535 1,302 458,453 99.57%
+Amazon 50,677 16,897 454,529 99.95%
+Gowalla 29,859 40,989 1,027,464 99.92%
 4.1. Datasets and Evaluation Metrics
-We conduct our experiments on two real-world datasets: MOOC and Anazon Video Games (Games for short).
+Table 3.
+Performance comparison of baselines and FourierKAN-GCF in terms of R@K and N@K.
+The superscript indicates the improvement is statistically significant where the p-value is less than 0.05.
+#T denotes seconds per epoch.
+Datasets MOOC Amazon Gowalla
+Metrics R@20 R@50 N@20 N@50 #T R@20 R@50 N@20 N@50 #T R@20 R@50 N@20 N@50 #T
+NGCF 0.3361 0.4799 0.1894 0.2349 5.2s 0.0379 0.0782 0.0196 0.0274 4.4s 0.1755 0.2811 0.1013 0.1270 37.5s
+LR-GCCF 0.3336 0.4809 0.1938 0.2294 5.1s 0.0440 0.0815 0.0224 0.0317 4.7s 0.1803 0.2971 0.1101 0.1369 40.1s
+LightGCN 0.3307 0.4773 0.1811 0.2217 4.3s 0.0447 0.0844 0.0227 0.0326 3.5s 0.1830 0.3044 0.1152 0.1414 26.4s
+UltraGCN 0.3194 0.4701 0.1962 0.2307 4.9s 0.0459 0.0844 0.0230 0.0331 4.0s 0.1798 0.2909 0.1059 0.1328 33.8s
+IMP-GCN 0.2788 0.4183 0.1717 0.2057 54.2s 0.0461 0.0839 0.0232 0.0323 44.9s 0.1808 0.2932 0.1060 0.1345 143.6s
+KAN-GCF 0.3417 0.4984 0.2024 0.2396 4.9s 0.0451 0.0837 0.0229 0.0325 4.0s 0.1922 0.3023 0.1142 0.1403 32.9s
+FourierKAN-GCF 0.3564 0.5065 0.2147 0.2462 4.6s 0.0473 0.0856 0.0252 0.0342 3.7s 0.1962 0.3077 0.1179 0.1436 29.8s
+w/o MD 0.3523 0.4912 0.2116 0.2449 4.5s 0.0452 0.0825 0.0221 0.0314 3.7s 0.1920 0.3031 0.1150 0.1408 29.5s
+w/o ND 0.3527 0.4839 0.2071 0.2439 4.5s 0.0452 0.0809 0.0219 0.0309 3.6s 0.1908 0.3017 0.1133 0.1389 29.6s
+Table 4. Performance comparison of different backbones on advanced recommendation models in terms of R@K and N@K.
+Datasets MOOC Amazon Gowalla
+Metrics Backbones R@20 R@50 N@20 N@50 R@20 R@50 N@20 N@50 R@20 R@50 N@20 N@50
+SimGCL LightGCN 0.3503 0.5032 0.2109 0.2428 0.0462 0.0837 0.0230 0.0327 0.2028 0.3126 0.1184 0.1459
+FourierKAN-GCF 0.3639 0.5188 0.2216 0.2543 0.0482 0.0868 0.0258 0.0352 0.2084 0.3163 0.1234 0.1521
+LightGCL LightGCN 0.2742 0.4139 0.1697 0.2018 0.0453 0.0818 0.0225 0.0318 0.1888 0.2976 0.1081 0.1348
+FourierKAN-GCF 0.2823 0.4291 0.1733 0.2072 0.0479 0.0864 0.0250 0.0347 0.2007 0.3100 0.1198 0.1452
+RecDCL LightGCN 0.3531 0.5009 0.2113 0.2423 0.0469 0.0848 0.0235 0.0334 0.1993 0.3052 0.1160 0.1433
+FourierKAN-GCF 0.3608 0.5120 0.2175 0.2514 0.0484 0.0863 0.0253 0.0349 0.2051 0.3113 0.1203 0.1472
+We conduct experiments on three real-world datasets: MOOC, Amazon Video Games (Amazon), and Gowalla.
 Details can be found in Table 2.
-For a fair comparison, we follow previous work (Zhou et al., 2023b) to sort all the observed user-item interactions of each dataset in chronological order based on the interaction timestamps.
+For a fair comparison, we sort all observed user-item interactions chronologically based on the interaction timestamps.
 Then, we split each dataset with a ratio of 7:1:2 for training, validation, and testing.
-Regarding evaluation metrics, we adopt two well-established metrics: Recall@K (R@K) and NDCG@K (N@K).
-We report the average metrics of all users in the test dataset among $K$ = 10, $K$ = 20, and $K$ = 50.
+Regarding evaluation metrics, we adopt two well-established metrics (Järvelin and Kekäläinen, 2002): Recall@K (R@K) and NDCG@K (N@K).
 4.2. Baselines and Experimental Settings
-To verify the effectiveness of FourierKAN-GCF, we select two traditional matrix factorization models (BPR-MF (Rendle et al., 2012) and BUIR (Lee et al., 2021)) and five state-of-the-art GCN-based models (NGCF (Wang et al., 2019), LR-GCCF (Chen et al., 2020), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), and IMP-GCN (Liu et al., 2021)) for comparison.
-For a fair comparison, we implement all baselines under a unified framework (Zhou et al., 2023b), fix the embedding size of both users and items to 64 for all models, initialize the embedding parameters with the Xavier initialization (Glorot and Bengio, 2010), and use Adam (Kingma and Ba, 2014) as the optimizer.
-Besides, we carefully tune the hyper-parameters of each baseline following their published papers to find optimal settings.
-For our FourierKAN-GCF, we tune the $\lambda$ of $L_{2}$ regularization term in {0, 1e-2, 1e-1, 1, 10}, and tune the layer number $L$ from 1 to 4.
-The gridsize $g$ of Fourier KAN is searched from {1, 2, 4, 8}.
-We further grid search the message dropout ratio $p_{m}$ and node dropout ratio $p_{n}$ in {0.0, 0.1, 0.2, 0.3}.
+To verify the effectiveness of FourierKAN-GCF, we select five GCN-based backbone models (NGCF (Wang et al., 2019), LR-GCCF (Chen et al., 2020), LightGCN (He et al., 2020), UltraGCN (Mao et al., 2021), and IMP-GCN (Liu et al., 2021)) as baselines.
+Moreover, FourierKAN-GCF can also be a more powerful backbone to enhance existing advanced self-supervised enhanced graph-based models (SimGCL (Yu et al., 2022), LightGCL (Cai et al., 2023), and RecDCL (Zhang et al., 2024)).
+For a fair comparison, we fix the embedding size of both users and items to 64 for all models, initialize embedding parameters with the Xavier initialization (Glorot and Bengio, 2010), and use Adam (Kingma and Ba, 2015) as optimizer.
+Besides, we tune the hyper-parameters of each baseline following their published papers.
+For FourierKAN-GCF, we fix the $\lambda$ = $1$ for $L_{2}$ regularization, and tune layer number $L$ from 1 to 4.
+The gridsize $g$ is searched from {1, 2, 4, 8}.
+Message and node dropout ratios $p_{m}$ and $p_{n}$ is searched from {0.0, 0.1, 0.2, 0.3}.
 4.3. Performance Comparison
 The results of our experiments are listed in Table 3.
-We can identify that the majority of GCN-based methods outperform traditional matrix factorization methods, which is a common trend.
-Our FourierKAN-GCF outperforms all baselines in two datasets across various metrics.
+Our FourierKAN-GCF outperforms all baselines in three datasets across various metrics.
+Moreover, FourierKAN-GCF is second only to LightGCN in efficiency, which demonstrates it effectively reduces the training difficulty associated with feature transformation.
 We owe our superiority to FourierKAN, which is easier to train and has greater representation power than MLP.
-It is worth noting that FourierKAN-GCF can adjust the training difficulty by adjusting the grid size $g$.
+Note that FourierKAN-GCF can adjust the training difficulty by adjusting the grid size $g$.
+4.4. Compatibility Analysis
+We further evaluate the compatibility of FourierKAN-GCF with advanced self-supervised graph-based models by replacing their backbones.
+As shown in Table 4, adopting FourierKAN-GCF significantly improves performance, highlighting the importance of feature transformation in GCF and establishing FourierKAN-GCF as an effective solution.
 4.5. Ablation Study
@@
 It shows that standard KAN is better than MLP but worse than our Fourier KAN.
+(a) Layer number $L$ (Recall@20)
+(b) Grid size $g$ (Recall@20)
+Figure 2. Effective of layer number $L$ and grid size $g$.
+(a) MOOC
+(b) Amazon
+(c) Gowalla
+Figure 3. Study of dropout ratio pair $p_{m}$ and $p_{n}$ (Recall@20).
 4.6. Sensitivity Analysis
-To analyze the hyper-parameter sensitivity of FourierKAN-GCF, we test the performance of FourierKAN-GCF on two datasets with different hyper-parameters.
-The optimal layer number $L$ is 3 on all datasets, shown in Figure LABEL:fig:l.
-Besides, Figure LABEL:fig:g demonstrates that for MOOC and Games datasets, optimal ($p_{m}$, $p_{n}$) pairs are (0.2, 0.2) and (0.1, 0.1), respectively.
-As Figure LABEL:fig:dropout illustrated, for the relatively dense dataset MOOC, $g$ = 4 is the best grid size.
-For the relatively sparse dataset Games, $g$ = 2 is the best grid size.
+To analyze the hyper-parameter sensitivity of FourierKAN-GCF, we test the performance of FourierKAN-GCF on three datasets with different hyper-parameters.
+The optimal layer number $L$ is 3 on all datasets, shown in Figure 2(a).
+Besides, Figure 2(b) demonstrates that for MOOC dataset, the optimal ($p_{m}$, $p_{n}$) pair is (0.2, 0.2), and for Amazon and Gowalla datasets, the optimal pairs are all (0.1, 0.1).
+As Figure 3 illustrated, for the relatively dense dataset MOOC, $g$ = 4 is the best grid size.
+For the relatively sparse datasets Amazon and Gowalla, $g$ = 2 is the best grid size.
 This further demonstrates that FourierKAN can be easily adapted to various datasets.
 5. Conclusion
-In this paper, we revisit the processes of feature transformation and nonlinear operations during message passing in the NGCF.
-We assert that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN but add complexity in training.
-To reduce the training difficulty while benefiting from these representational enhancements, we introduce a novel feature transformation method called FourierKAN.
-Inspired by KAN, FourierKAN employs a Fourier Coefficients instead of the Spline function in standard KAN, which further reduces the difficulty of training while increasing the representation power.
-Furthermore, based on FourierKAN, we introduce a simple and effective graph-based recommendation model (FourierKAN-GCF) with extra dropout strategies.
-Extensive experiments on public datasets demonstrate the superiority of our model over the state-of-the-art methods.
+In this paper, we revisit feature transformation and nonlinear operations in the message-passing mechanism of GCNs.
+While feature transformation enhances interaction representation and boosts performance, it increases training complexity.
+To this end, we introduce a new feature transformation for GCF called FourierKAN.
+Inspired by KAN, FourierKAN employs the Fourier Coefficients instead of the Spline function in standard KAN.
+We further propose a simple yet effective GCF model (FourierKAN-GCF), which reduces the difficulty of training.
+In addition, FourierKAN-GCF can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance.
+Extensive experiments on public datasets verify the superiority of our model over the advanced methods.
+GenAI Usage Disclosure
+No GenAI tools were used in any stage of the research, nor in the writing.
+Acknowledgements.
+This work was supported by the Hong Kong UGC General Research Fund no. 17203320 and 17209822, and the project grants from the HKU-SCF FinTech Academy.